OpenStudy (anonymous):

What is the final balance for the investment? $1500 for 3 years at 10% compounded annually A. $1875.00 B. $1625.00 C. $1996.50 D. $2110.50

OpenStudy (calculusxy):

Remember the exponential function \(f(x) = a(b)^t\) a = principal amount b = rate t = time

OpenStudy (calculusxy):

Starting (principal) amount is what?

OpenStudy (anonymous):

1500?

OpenStudy (calculusxy):

Correct. The rate?

OpenStudy (anonymous):

3

OpenStudy (calculusxy):

Nope. That's the time actually because it's 3 years.

OpenStudy (anonymous):

oh so .010?

OpenStudy (calculusxy):

So for the rate, what actually happens is that you add 1 to the decimal form of the percent (10%).

OpenStudy (anonymous):

1.10?

OpenStudy (anonymous):

im confused and have to go

OpenStudy (calculusxy):

Yes that's correct for the rate. And the time you said already was 3 years.

OpenStudy (calculusxy):

So substitute the corresponding values into their respective places in the formula \[f(n) = 1500(1.1)^3\] Evaluate that.

OpenStudy (anonymous):

umm lets see

OpenStudy (anonymous):

=C

OpenStudy (calculusxy):

Correct!

OpenStudy (anonymous):

ty so much ill be back on in a little bit